### v2

# Improving generalisation via anchor multivariate analysis

Abstract
We introduce a causal regularisation extension to anchor regression (AR) for improved out-of-distribution (OOD) generalisation. We present anchor-compatible losses, aligning with the anchor framework to ensure robustness against distribution shifts. Various multivariate analysis (MVA) algorithms, such as (Orthonormalized) PLS, RRR, and MLR, fall within the anchor framework. We observe that simple regularisation enhances robustness in OOD settings. Estimators for selected algorithms are provided, showcasing consistency and efficacy in synthetic and real-world climate science problems. The empirical validation highlights the versatility of anchor regularisation, emphasizing its compatibility with MVA approaches and its role in enhancing replicability while guarding against distribution shifts. The extended AR framework advances causal inference methodologies, addressing the need for reliable OOD generalisation.
Keywords Distribution shift  $\cdot$ Robust estimation  $\cdot$ Domain generalisation  $\cdot$ Causality  $\cdot$ Multivariate analysis  $\cdot$ Detection and attribution  $\cdot$ Climate Change
1 Introduction
Data sources in contemporary machine learning applications are often heterogeneous, leading to potential distribution shifts Sugiyama and Kawanabe (2012); Shen et al. (2021). This is a particularly relevant problem in computer vision Csurka (2017), healthcare Zhang et al. (2021), finance, Earth and climate sciences (Tuia et al., 2016; Kellenberger et al., 2021) and social sciences, as variations in data patterns can significantly impact model performance and generalisation in the out-of-distribution (OOD) setting, also referred to as domain generalisation (Shen et al., 2021; Zhou et al., 2023). Various frameworks have been proposed to formally address the emergence of distribution shifts during the testing phase Peters et al. (2016); Arjovsky et al. (2020). In cases where the data distribution is entailed by a Structural Causal Model (SCM) Peters et al. (2017), one can consider distribution shifts arising from intervention on specific variables of the SCM. Notably, the Instrumental Variable (IV) regression exhibits robustness to arbitrarily strong interventions (Bowden and Turkington, 1990).
However, pursuing algorithms robust to strong interventions may be overly conservative, especially when prior knowledge is available regarding the intervention strength that generates the distribution shift. Anchor Regression (AR) addresses this challenge by explicitly considering interventions on exogenous variables up to a specified strength (Rothenhäusler et al., 2018). This approach results in a straightforward and manageable regularisation of the classical Ordinary Least Squares (OLS) method, offering a solution for achieving distributionally robust linear regression.
Several AR extensions have been proposed, demonstrating its versatility in handling different scenarios. Notable examples include (1) Kernel Anchor Regression (Shi and Xu, 2023), which extends AR to nonlinear SCMs by working on Reproducing Kernel Hilbert Spaces; (2) Targeted Anchor Regression Oberst et al. (2021), which exploits prior knowledge of the direction of intervention on the anchor variables; and (3) Proxy Anchor Regression Oberst et al. (2021), which addresses cases where the anchor is unobserved but proxy variables are available. Closer to our work, Kook et al. (2022) introduce a distributional extension of AR, expanding its application beyond the $\ell_{2}$ loss to accommodate censored or discrete targets.
All the literature above focuses on regression problems with a one-dimensional target variable and mainly employs least-squared error as the loss function. This paper extends the anchor framework introduced by Rothenhäusler et al. (2018) and contributes in two distinct ways. We first broaden its applicability to various loss functions, ensuring compatibility and robustness against distribution shifts caused by constrained interventions on the anchor variables. This extension redefines conventional MVA algoritms (Borchani et al., 2015; Arenas-Garcia et al., 2013), within the anchor framework, including algorithms such as Partial Least Squares (PLS), Orthonormalized PLS or Reduced Rank Regression (RRR) and Multilinear Regression (MLR). This signifies that their respective loss functions are anchor-compatible , and applying straightforward regularisation enhances their robustness in testing.
After reviewing the AR framework in §2, we study robustness through AR-based regularisation in §3 (anchor-compatible losses and their role in MVAs) and discuss the key links to causality and invariance (cf. §4). Then, we provide estimators for the selected algorithms and examine their consistency (cf. Section 5), supporting our theoretical findings through an extended set of simulation experiments. Finally, we showcase their efficacy in a challenging real-world climate science problem (cf. §6.2). We conclude in §7. The presented AR framework enhances causal inference methods by improving robust OOD generalisation. The code used in this work can be found at this GitHub repository.
2 Anchor framework
This section delineates the class of linear SCMs relevant to our work (§2.2) and offers insight into why anchor regularisation can be employed to obtain distributionally robust estimators (§2.3) for a more general class than the $\ell_{2}$ loss.
2.1 Notation
Bold capital letters $\mathbf{U}$ represent matrices, bold lowercase letters $\mathbf{u}$ denote vectors, regular lowercase letters $u$ signify scalars, and uppercase letters $U$ represent random variables (potentially multivariate) which for notation sakes we assume centred. Given two random vector $U$ and $V$ we define $UV$ as the row-wise stacking operation $UV=(U^{T},V^{T})^{T}$. The symbol $C_{UV}$ denotes Kronecker products involving random variables, i.e. $C_{UV}=UV\otimes UV$and to simplify the notation we write $C_{U}=U\otimes U$. Since we assume centred random variables $\Sigma_{UV}=\mathbb{E}[C_{UV}]$ and $\Sigma_{U}=\mathbb{E}[C_{U}]$. By a slight abuse of notation, we denote $C_{UV|A}=\mathbb{E}\left[UV|A\right]\otimes\mathbb{E}\left[UV|A\right]$ and $\Sigma_{UV|A}$ its expectation. [^1] Not to be confused with the conditional covariance matrix $Cov(UV|A)$. Our notation statisfies the total variance-covariance formula $\Sigma_{UV}=\mathbb{E}[Cov(UV|A)]+\Sigma_{UV|A}$. Again for simplification, we write $\sigma_{U}=\mathbb{E}[C_{U}]$ and $\sigma_{UV}=\mathbb{E}[U\otimes V]$. We denote the identity matrix with $\mathbf{I}$, and its dimension is implied by the context when not explicitly stated. Finally, .$\|.\|_{F}$ represents the Froebinius norm.
2.2 The Class of Anchor SCM
$X$$Y$$H$$A$
Figure 1: DAG studied in this work and induced by the SCM in (7).
In the following, we assume the Directed Acyclic Graph (DAG) in Fig. 1, where $X\in\mathbb{R}^{d}$ denotes the observed covariates, $Y\in\mathbb{R}^{p}$ represents the response (or target) variables, $H\in\mathbb{R}^{r}$ are unobserved variables potentially confounding the causal relation between $X$ and $Y$ and $A\in\mathbb{R}^{q}$ the exogenous (so-called anchor) variables. Following Rothenhäusler et al. (2018), we assume that the distribution of $(X,Y,H)$ is entailed by the following linear SCM $\mathcal{C}$:
$\begin{pmatrix}X\\ Y\\ H\end{pmatrix}=\mathbf{B}\begin{pmatrix}X\\ Y\\ H\end{pmatrix}+\varepsilon+\mathbf{M}A,$ (7)
where $\mathbf{B}\in\mathbb{R}^{(d+p+r)\times(d+p+r)}$ and $\mathbf{M}\in\mathbb{R}^{(d+p+r)\times q}$ are unknown constant matrices and $\varepsilon\in\mathbb{R}^{d+p+r}$ is a vector of random noise. We assume that $A$ and $\varepsilon$ are independent and have finite variances, $\varepsilon$ components are independent, and $X$ and $Y$ have zero mean. Now, assuming that $(\mathbf{I}-\mathbf{B})$ is invertible, Eq. (7) can be written as:
$\begin{pmatrix}X\\ Y\\ H\end{pmatrix}=(\mathbf{I}-\mathbf{B})^{-1}(\varepsilon+\mathbf{M}A).$ (8)
The existence of $(\mathbf{I}-\mathbf{B})^{-1}$ is satisfied if the linear SCM is acyclic. Since our main interest lies in defining the variance-covariance $\Sigma_{XY}$ of $X$ and $Y$, we use $\mathbf{D}\in\mathbb{R}^{(d+p)\times(d+p+r)}$ to express:
$\begin{pmatrix}X\\ Y\end{pmatrix}=\mathbf{D}(\varepsilon+\mathbf{M}A),$ (9)
whose existence is implied by the existence of $(\mathbf{I}-\mathbf{B})$ inverse.
2.3 Anchor-Optimal Space
Let us now build upon (9) to illustrate how distributional robustness can be formulated in a causal context and give some intuition of the extension we propose in this work. Later, we will delve into the crucial role of $\Sigma_{XY}$ and its decomposition with $\varepsilon$ and anchor variables $A$. We focus on creating algorithms utilising $\Sigma_{XY}$. Note that, from (9), we can easily express $\Sigma_{XY}$ as
$\Sigma_{XY}$ $=\mathbf{D}\Sigma_{\varepsilon}\mathbf{D}^{T}+\mathbf{D}\mathbf{M}\Sigma_{A}\mathbf{M}^{T}\mathbf{D}^{T},$ (10)
which combines the variance from observed exogenous (anchors) and unobserved exogenous (noise) variables.
Let us consider the triplet $(X,Y,H)\sim\operatorname{\mathbb{P}_{\nu}}$, where $\operatorname{\mathbb{P}_{\nu}}=\mathbb{P}^{do(A:=\nu)}_{\mathcal{C}}$ represents the perturbed distribution resulting from intervention (Peters et al., 2017) on the anchor variable $A$, where the distribution of $\nu$ is assumed to be mean-centred. From (10), the interventional variance-covariance matrix $\Sigma_{XY}^{do(A:=\nu)}$ can be expressed as
$\Sigma_{XY}^{do(A:=\nu)}$ $=\mathbf{D}\Sigma_{\varepsilon}\mathbf{D}^{T}+\mathbf{D}\mathbf{M}\Sigma_{\nu}\mathbf{M}^{T}\mathbf{D}^{T}.$ (11)
By imposing the constraint $\Sigma_{\nu}\preceq\gamma\Sigma_{A}$ [^2] Here, $U\preceq V$ means that $V-U$ is positive semi-definite. (the rationale for which will be developed in section §3.1), we can bound the interventional variance-covariance matrix $\Sigma_{XY}^{do(A:=\nu)}$ using the training covariances $\Sigma_{\varepsilon}$ and $\Sigma_{A}$. The anchor regularizer parameter $\gamma\in\mathbb{R}^{+}$, controls the amount of causal regularisation by constraining interventions on $A$. This yields the specific distribution class:
$C^{\gamma}=\{\operatorname{\mathbb{P}_{\nu}}:\Sigma_{\nu}\preceq\gamma\Sigma_{A}\}.$ (12)
As demonstrated in Eq. (31) and Eq. (30), both $\Sigma_{\varepsilon}$ and $\Sigma_{A}$ can be expressed in terms of $\Sigma_{XY|A}$ and $\Sigma_{XY}$ of the training distribution, which leads to the following inequality:
$\Sigma_{XY}^{do(A:=\nu)}\preceq\Sigma_{XY}+(\gamma-1)\Sigma_{XY|A}.$ (13)
Consequently, the interventional covariance matrix $\Sigma_{XY}^{do(A:=\nu)}$
can be bounded using only training data according to Eq. (13). As we will develop in Section 3.1, all algorithms exploiting $\Sigma_{XY}$ can thus be anchor-regularised, leading to distributionally robust estimators.
An example is the $\ell_{2}$ loss, defined for a one-dimensional response variable as $\mathbb{E}[\mathcal{L}(X,Y;\mathbf{b})]=\mathbb{E}[(Y-\mathbf{b}^{T}X)^{2}]=\Sigma_{Y}-2\mathbf{b}^{T}\sigma_{XY}+\mathbf{b}^{T}\Sigma_{X}\mathbf{b}$, which is evidently linear on $\Sigma_{XY}$. Similarly, the PLS algorithm Abdi (2010) aims to find a pair of projection matrices $\mathbf{W}_{x}\in\mathbb{R}^{d\times u}$ and $\mathbf{W}_{y}\in\mathbb{R}^{p\times v}$ that maximise the covariance between the transformed data $X\mathbf{W}_{x}$ and $Y\mathbf{W}_{y}$, i.e. assuming centered $X$ and $Y$, we aim to maximise $\text{tr}(\mathbf{W}_{x}^{T}\Sigma_{XY}\mathbf{W}_{y})$ while constraining $\mathbf{W}_{x}$ and $\mathbf{W}_{y}$ to be orthogonal. Here $u$ and $v$ represent the dimension of the projected spaces. In this case, as the loss is also linear on matrix $\Sigma_{XY}$, i.e. PLS can be anchor-regularised, cf. §3.1.
An intuitive way of thinking about anchor regularisation is to consider that by perturbing the training data (see Eq. (21)), one obtains a dataset that is close to the worst-case scenario we could observe during testing. These observations leads us to consistently define a class of loss functions and demonstrate their distributional robustness properties over the class of perturbed distributions $C^{\gamma}$ defined in (12).
3 Robustness through anchor regularisation
Given $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$, distributionally robust estimation can be formulated as finding a minimiser of the supremum of the population loss under a certain class of distributions:
$\operatorname{\text{argmin}}_{\mathbf{\Theta}}\sup_{\mathcal{Q}\in\mathbb{Q}}\mathbb{E}_{(X,Y)\sim\mathcal{Q}}[\mathcal{L}(X,Y;\mathbf{\Theta})],$ (14)
where $\mathbb{Q}$ is a class of distributions and $\mathcal{L}(X,Y;\mathbf{\Theta})$ is a loss function over $X$ and $Y$ with parameters $\mathbf{\Theta}$. In essence, we aim for our model to perform well even in the worst-case scenario when the distribution of $(X,Y)$ belongs to the class of $\mathbb{Q}$. The choice of the class $\mathbb{Q}$ is particularly important as different definitions lead to various types of robustness.
The class $\mathbb{Q}$, often referred to as the uncertainty set, can be constrained using methods such as $f$-divergence (see, e.g., Namkoong and Duchi, 2016) or the Wasserstein distance (see, e.g., Esfahani and Kuhn, 2015). The challenge of OOD generalisation is also closely related to the field of transfer learning (see, e.g., Zhuang et al., 2021). For an in-depth review of OOD generalisation, we recommend consulting (Shen et al., 2021; Zhou et al., 2023).
Of particular interest for this work, causal inference could be understood as a distributionally robust estimation where the class $\mathbb{Q}$ is the class of all distributions arising from arbitrarily large interventions on the variables of an SCM (see Shen et al., 2021, section 4.1.1).
Besides, Rothenhäusler et al. (2018) proposed to bound the intervention strength on the exogenous variable and relax the assumption that it only affects $Y$ through $X$, considering the linear SCM in Fig. 1. By considering the set of distributions $C^{\gamma}$ defined in (12), they demonstrate that their estimator is optimal in the sense of (14) for the $\ell_{2}$ loss.
3.1 anchor-compatible loss
The distributional robustness properties stemming from intervention on the anchor variables arise from the ability to bound the $\Sigma^{do(A:=\nu)}_{XY}$ using the observational covariance matrix $\Sigma_{XY}$ and $\Sigma_{XY|A}$. Consequently, any loss function expressed as a linear function of $\Sigma_{XY}$ can attain distributional robustness similar to AR using anchor regularisation.
Definition 3.1 (anchor-compatible loss).
We say that a loss function $\mathcal{L}(X,Y;\mathbf{\Theta})$ is anchor-compatible if it can be written as
$\mathcal{L}(X,Y;\mathbf{\Theta})=f_{\mathbf{\Theta}}(C_{XY}),$ (15)
where $f_{\mathbf{\Theta}}:\mathbb{R}^{d\times p}\to\mathbb{R}$ is a linear form and $\mathbf{\Theta}$ are the parameters to be learned.
This theorem extends Th. 1 in Rothenhäusler et al. (2018) to the previously defined class of loss functions.
Theorem 3.2.
Let the distribution of $(X,Y,H)$ be entailed by the SCM (7) and $\mathcal{L}(X,Y;\mathbf{\Theta})$ be an anchor-compatible loss function. Then for any set of parameters $\mathbf{\Theta}$, and any causal regulariser $\gamma\in\mathbb{R}^{+}$, we have
$\begin{split}\sup_{\nu\in C^{\gamma}}&\operatorname{\mathbb{E}_{\nu}}[\mathcal{L}(X,Y;\mathbf{\Theta})]=f_{\mathbf{\Theta}}(\Sigma_{XY})\\ +&(\gamma-1)f_{\mathbf{\Theta}}(\Sigma_{XY|A}),\end{split}$ (16)
where $C^{\gamma}=\{\operatorname{\mathbb{P}_{\nu}}:\Sigma_{\nu}\preceq\gamma\Sigma_{A}\}$.
Proof can be found in the appendix (3.2).
Let us note that the term $f_{\mathbf{\Theta}}(\Sigma_{XY|A})$ could be understood as a causal regulariser enforcing invariance of the loss w.r.t the anchor variable.
For $\gamma=1$, anchor regularisation aligns with unregularised estimation. The theorem suggests that in the case of distributional shifts observed during testing due to constrained intervention on $A$, the anchor-regularised estimator
$\mathbf{\Theta}^{\gamma}=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\Sigma_{XY})+(\gamma-1)f_{\mathbf{\Theta}}(\Sigma_{XY|A})$ (17)
acts as optimal estimators, as per the definition provided in Eq. (14). We will proceed to demonstrate the practical application of this general result to standard Multivariate Analysis (MVA) algorithms.
3.2 Common Multivariate Analysis Algorithms
Table 1: Characterisation of common MVAs with anchor regularisation (loss, constraints, and anchor-compatibility).
MLR OPLS RRR PLS CCA
Loss $\|Y-\mathbf{W}^{T}X\|_{F}^{2}$ $\|Y-\mathbf{U}\mathbf{V}^{T}X\|_{F}^{2}$ $\|Y-\mathbf{W}X\|_{F}^{2}$ $\text{tr}\left(\mathbf{W}_{x}^{T}X^{T}Y\mathbf{W}_{y}\right)$ $\text{tr}\left(\mathbf{W}_{x}^{T}X^{T}Y\mathbf{W}_{y}\right)$
[…]
Comp.
A common and straightforward approach to extend least squares (LS) regression to multivariate settings is treating responses independently in multioutput regression (Borchani et al., 2015). This approach aims to identify the regression coefficients $\mathbf{W}$ such that $\hat{Y}=\mathbf{W}^{T}X$, e.g. by solving the LS problem $\widehat{\mathbf{W}}=\operatorname{\text{argmin}}_{\mathbf{W}}\|Y-\mathbf{W}^{T}X\|^{2}_{F}$. When the columns of $Y$ exhibit correlation, RRR addresses a similar optimisation problem assuming that the rank of $\mathbf{W}$ is lower than $\min(d,p)$, see Izenman (1975). Alternatively, in OPLS Arenas-García and Gómez-Verdejo (2015), we do not assume that $\mathbf{W}$ is low rank. Instead, $\mathbf{W}$ is determined by imposing the constraints $\mathbf{W}=\mathbf{V}^{T}\mathbf{U}$, where both $\mathbf{U}$ and $\mathbf{V}$ have a rank $\rho\leq\min(d,p)$.
Additionally, OPLS is commonly solved through matrix deflation, eigenvalue, or generalised eigenvalue decomposition Arenas-García and Gómez-Verdejo (2015). An advantageous property of OPLS is that the columns of the solution $\mathbf{V}$ are ordered based on their predictive performance on $Y$. This is valuable when predicting $Y$ is not the sole objective, and there is interest in learning a relevant information-retaining subspace of $X$ (here $\mathbf{V}^{T}X$). These are natural extensions of LS regression and are anchor-compatible. Please refer to Appendix 8 for further details and proofs.
PLS and CCA are typically preferred when prioritizing learning a pertinent subspace of $X$ over predictive performance. These algoritms aim to maximise the similarity between the latent representations of predictor and target variables. Specifically, PLS assumes the following latent representation:
$X$ $=\mathbf{P}^{T}T+N_{x},$
$Y$ $=\mathbf{Q}^{T}U+N_{y},$
and seeks to maximise the covariance between $\mathbf{W}_{x}^{T}X$ and $\mathbf{W}_{y}^{T}Y$, while CCA aims to maximise the correlation between the estimated latent spaces. Correlation as a similarity measure ensures equal importance for each dimension in the learned latent space, independent of data variance.
PLS, being anchor (see Proposition 8.2), benefits from causal regularisation for distributional robustness. In contrast, CCA lacks anchor compatibility due to nonlinearity in its loss function concerning variance-covariance (see §8.3). As illustrated in Fig. 2, CCA shows reduced robustness under interventions on the anchor variable’s variance. Recognizing the established equivalence between CCA and OPLS (Sun et al., 2009), it may appear puzzling that one algorithm is anchor-compatible while the other is not. We delve deeper into this matter in the supplementary (§8).
4 Towards causality, invariance and stability
Given the well-established connection between causality and invariant predictions (see, for instance Bühlmann (2020) Prop. 3.1), exploring the relationship between anchor regularisation, causality, and invariance could offer valuable insights.
For instance, IV regression assumes that the noise term $\varepsilon_{y}$ is independent of the instrument (or anchor) variable $A$ (i.e., $\varepsilon_{y}\perp\!\!\!\perp A)$ thus IV estimations can be performed using the moment restriction Newey and Powell (2003) $\mathbb{E}[Y-b^{T}X|A]=0$, which can be related to the right term of Eq. (16) for anchor-regularised $\ell_{2}$ loss. [^3] The right term is exactly $\mathbb{E}[(\mathbb{E}[Y-b^{T}X|A])^{2}]$ in AR. Thus, as $\gamma\to\infty$, the anchor-regularised loss converges to the IV estimation, i.e.
$\mathbf{\Theta}^{IV}$ $=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\Sigma_{XY|A})=\lim_{\gamma\to+\infty}\mathbf{\Theta}^{\gamma}.$ (18)
Increasing $\gamma$ ensures stable predictive performance by promoting robustness to shifts in the distribution of $X$ resulting from interventions on the anchor.
Another notable scenario is the partialling-out (PA) setting, closely related to double machine learning (see e.g. Chernozhukov et al., 2017), where the anchor variables directly influence both $X$ and $Y$, and there is no hidden confounder $H$.
In this setup, the causal effect is estimated by deriving residuals of $X$ and $Y$ predicted by $A$, followed by estimating the regression coefficient of the residuals of $X$ on the residuals of $Y$. This process is equivalent to estimating the causal effect of $X$ on $Y$ within the orthogonal complement space of $A$. Thus, an anchor-regularised algorithm with $\gamma=0$ is akin to PA estimation.
$\mathbf{\Theta}^{PA}$ $=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\Sigma_{XY|A^{\perp}})=\lim_{\gamma\to 0}\mathbf{\Theta}^{\gamma},$ (19)
where $A^{\perp}=\mathbf{I}-A$ is the orthogonal complement of $A$. Hence, PA solves the minimax problem Eq. (16) for an empty set Eq. (12) and is thus optimal under distribution shift arising from null intervention on $A$. Note that, in some instances, the parameters $\mathbf{\Theta}^{\gamma}$ remain consistent across all $\gamma\in[0,\infty]$, resulting in what we term as anchor stability. Under these circumstances, and subject to relatively mild assumptions outlined in Rothenhäusler et al. (2018) section 3.2, the parameters $\mathbf{\Theta}$ can be causally interpreted, as demonstrated in (Rothenhäusler et al., 2018, Th. 4), indicating the absence of a confounder $H$ between $X$ and $Y$.
5 Estimators
In the previous sections, we derived properties of various MVA methods in the population case. We consider now the sample setting where we are given $n$ i.i.d observations of $(X,Y,A)$. As customary, observations are collectively organised row-wise, in the following matrices, $\mathbf{X}\in\mathbb{R}^{n\times d}$, $\mathbf{Y}\in\mathbb{R}^{n\times p}$ and $\mathbf{A}\in\mathbb{R}^{n\times q}$. For a given anchor-compatible loss $\mathcal{L}(X,Y;\mathbf{\Theta})=f_{\mathbf{\Theta}}(C_{XY})$, we propose to consider a simple estimator of the population parameter $\mathbf{\Theta}^{\gamma}$, obtained by the plug-in principle:
$\widehat{\mathbf{\Theta}}^{\gamma}=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\mathbf{S}_{XY})+(\gamma-1)f_{\mathbf{\Theta}}(\mathbf{S}_{XY|A}),$ (20)
where $\mathbf{S}_{XY}$ and $\mathbf{S}_{XY|A}$ are respectively the empirical estimators of $\Sigma_{XY}$ and of $\Sigma_{XY|A}$.
As proved by Rothenhäusler et al. (2018), $\widehat{\mathbf{\Theta}}^{\gamma}$ can be computed by transforming the training data: [^4] Proof of the equivalence of using transformed data as in Eq. (21) is equivalent to Eq. (20) is available in Supplementary  9.2. From Rothenhäusler et al. (2018), the transformed data set $(\tilde{\mathbf{X}},\tilde{\mathbf{Y}})$ can be interpreted as an artificially perturbed data.
$\begin{split}\tilde{\mathbf{X}}&=(\mathbf{I}+(\sqrt{\gamma}-1)\Pi_{\mathbf{A}})\mathbf{X},\\ \tilde{\mathbf{Y}}&=(\mathbf{I}+(\sqrt{\gamma}-1)\Pi_{\mathbf{A}})\mathbf{Y},\end{split}$ (21)
where $\Pi_{\mathbf{A}}=\mathbf{A}(\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}^{T}\in\mathbb{R}^{n\times n}$ projects onto the column space of $\mathbf{A}$, assuming that $\mathbf{A}^{T}\mathbf{A}$ is invertible and $\mathbf{X}$ and $\mathbf{Y}$ are centered.
Finally, build the estimator $\widehat{\mathbf{\Theta}}^{\gamma}=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\mathbf{S}_{\tilde{X}\tilde{Y}})$, where $\mathbf{S}_{\tilde{X}\tilde{Y}}$ is the empirical variance-covariance of the projected data defined in Eq.  (21).
Consistency
From the law of large numbers, the empirical covariance matrix of $(X,Y,A)$ converges to its empirical covariance matrix, i.e., $\mathbf{S}_{XY}\xrightarrow[n\to\infty]{}\Sigma_{XY}$. Thus, the consistency of the anchor-regularised estimator depends on the consistency of the original estimator. For continuous functions $f$, we have by continuity that $\mathbf{\widehat{\Theta}}^{\gamma}\xrightarrow[n\to\infty]{}\mathbf{\Theta}^{\gamma}$ (see Rothenhäusler et al., 2018, section 4.1).
Computational Complexity
The computational cost of projecting $X$ and $Y$ into the span of $A$ involves computing the covariance matrix of $A$ (of cost $O(nr^{2})$), inverting it (of cost $O(r^{3})$), and two matrix products (of cost $O(nrp)$ and $O(nr^{2})$), resulting in a total complexity of $O(nr^{2}+nrp+r^{3})$. Assuming the unregularised MVA algorithm’s computational cost is $c$, the complexity of its anchor-regularised version is $O(c+nr^{2}+nrp+r^{3})$.
Parameter selection
When knowledge about the graph and anchor variable distributional shift is available, one can estimate the optimal $\gamma$ for robustness against worst-case scenarios. For instance, if the expected anchor variable perturbation strength is up to $1.5$, Rothenhäusler et al. (2018) recommends $\gamma=1.5$. Refer to Section 8.2 of Rothenhäusler et al. (2018) for a detailed estimation approach. Without prior knowledge, various strategies can be considered: Rothenhäusler et al. (2018) propose $\gamma=2$ as a suitable default choice, while Sippel et al. (2021) and Székely et al. (2022) suggest selecting $\gamma$ as a trade-off between prediction error (MSE or $R^{2}$ score) and the correlation between residuals and the anchor variable, or the value of projected residuals in the anchor variable’s span. Similar strategies have been proposed in deemed related algorithms Cortés-Andrés et al. (2022); Li et al. (2022).
Figure 2: Robustness to increasing perturbation strength for PA, unregularised, IV, anchor-regularised ($\gamma=5$) algorithms ( MLR, OPLS, RRR, PLS and CCA). Anchor versions show robustness in terms of $R^{2}$ for bounded intervention strength.
High Dimensional Estimators
Dealing with high dimensional data challenges estimators and often requires introducing regularisation Bühlmann and Van De Geer (2011); Candes and Tao (2007). Regularisation has been generally introduced to control models’ capacity and avoid overfitting Girosi et al. (1995); Tibshirani (1996); Hastie et al. (2009), but also to introduce prior (domain) knowledge in the algorithms, as in AR. Adding a regularisation term $\Omega(\|\mathbf{\Theta}\|)$ to the empirical loss in Eq. (20) leads to estimate $\mathbf{\Theta}$ as
$\begin{split}\widehat{\mathbf{\Theta}}^{\gamma}=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}&(\mathbf{S}_{XY})\\ +(\gamma-1)&f_{\mathbf{\Theta}}(\mathbf{S}_{XY|A})+\Omega(\|\mathbf{\Theta}\|).\end{split}$ (22)
In CCA, OPLS, and PLS regression (or the rank in RRR), the number of components can be viewed as regularisation hyperparameters. Consequently, the optimisation task for anchor-regularised MVA algorithms may involve many hyperparameters. For instance, in anchor-regularised Reduced Rank Ridge Regression (A-RRRR), three hyperparameters require tuning: $\ell_{2}$ regularisation $\alpha$, rank $\rho$, and anchor regularisation $\gamma$. Since each hyperparameter optimization addresses different objectives, it may be impractical to optimise all of them simultaneously. Sippel et al. (2021) advise selecting hyperparameters as a trade-off between prediction performance (e.g., MSE or $R^{2}$) and a proxy of robustness to anchor intervention (e.g. correlation between anchor and residuals).
6 Experiments
Our theoretical findings are substantiated through an extensive series of experiments that highlight the robustness of anchor-regularised MVA algorithms to increasing perturbation strength (cf. Section 6.1). Furthermore, we provide a concrete high-dimensional example to elucidate the process of hyperparameter selection (cf. Section 6.1), and demonstrate its practical relevance by showcasing how anchor regularisation enhances climate predictions in a real-world Climate Science application (cf. Section 6.2).
6.1 Simulation Experiments
Robustness to Perturbation Strength
To demonstrate how anchor regularisation can be applied with different multivariate analysis algorithms, we adapt the experiments from (Rothenhäusler et al., 2018, Sec. 2.2) to a multioutput setting.
We assume that the training data $(A,X,Y)$ follows a distribution based on the linear SCM:
$\begin{split}\varepsilon_{A},\varepsilon_{H},&\varepsilon_{X},\varepsilon_{Y}\sim\mathcal{N}(0,1)\\ A&\leftarrow\varepsilon_{A}\\ H&\leftarrow\varepsilon_{H}\\ X&\leftarrow A+H+\varepsilon_{X}\\ Y&\leftarrow\mathbf{W}^{T}X+H+\varepsilon_{Y},\end{split}$ (23)
where $X$ and $Y$ are multivariate, of respective dimension $d=p=10$ ($d=p=300$ with $n=200$ in the high dimensional setting), and $\mathbf{C}$ is low rank (see Supplementary  10.1 for more details on the experiments settings). We generate test data by intervening on the anchor variable’s distribution, setting $A\sim\mathcal{N}(0,t)$, where $t$ is the perturbation strength. Each MVA algorithm assumes oracle knowledge of the rank $\rho$ of $\mathbf{C}$, aligning RRR’s rank, PLS regression’s component count, and CCA accordingly. IV setting ($A$ only affects $Y$ through $X$) and we show that even in this scenario, anchor regularised algorithms (with $\gamma=5$) outperform the PA, IV, and unregularised algorithms.
Shown in Fig. 2, all anchor-compatible algorithms exhibit robustness to distribution shifts. Anchor-regularised models (with $\gamma=5$) excel with bounded-strength interventions; PA regularisation is optimal for weak interventions; and IV regularisation for unlimited perturbation strength. Overall, anchor-regularised algorithms maintain stable performance across various perturbation strengths.
We also conducted a set of high-dimensional experiments ($p\gg n$ and $d\gg n$) using RRR and MLR both regularised with $\ell_{2}$ norm (ridge regularisation). In both cases, the anchor-regularised algorithms exhibit robustness to increasing perturbation strength (see Fig. 6 in §10.1). This is also the case when $A$ is a confounder (affecting both $X$ and $Y$, potentially through $H$), in which case anchor-regularised algorithms present are equally optimal for a wide range of perturbation strength (see Fig. 8 in §10.1). Though not as pronounced as in anchor-compatible algorithms, anchor-regularised CCA also displays robustness to perturbations in $A$, prompting further investigation into its behaviour.
Hyperparameter Selection
We showcase how hyperparameter selection can be performed through a simulation experiment in a high-dimensional setting using an anchor-regularised version of Reduced Rank Ridge Regression (Mukherjee and Zhu, 2011).
In this setting, we are given three hyperparameters: $\gamma$ enforces robustness to intervention on the anchor, $\alpha$ and $\rho$ aim to maximise prediction performance. Following (Sippel et al., 2021) we select hyperparameters as a trade-off between predictive performance (measured via MSE) and correlation between anchor and residuals. We give equal weights to both objectives but this can be adapted regarding the knowledge available and the application. Thus hyperparameters are selected at testing such that they minimise the combination of the two objectives. Fig. 3 shows how augmenting regularisation for distributional robustness to anchor intervention often reduces predictive performance in the training sample (Fig. 3.A), yet the trade-off translates into improved predictive performance in testing samples. More details about this experiment setting are in  §10.2.
Figure 3: Pareto front for hyperparameters $(\gamma,\alpha,\rho)$ in RRRR and A-RRRR. For each pair $(\gamma,\alpha)$, $\rho$ is selected to minimise the weighted sum of objectives (equal weights given here).
6.2 Real Experiment: Robust Climate Prediction
We showcase the efficacy of our approach in a real-world application within the Detection and Attribution of Climate Change (D&A) domain. We extend the methodology of Sippel et al. (2021), who utilised AR for robust detection of forced warming against increased climate variability, by applying it to predict multidimensional local climate responses. Given the increased variability observed in recent climate models Parsons et al. (2020) and observations (DelSole, 2006; Kociuba and Power, 2015; Cheung et al., 2017), our approach aims to ensure robustness against potential underestimation of decadal and multidecadal internal climate variability (Parsons et al., 2020; Mcgregor et al., 2018).
Objective
Specifically, our objective is to use a $p$-dimensional temperature field $X_{\text{mod}}$ to predict a temperature response $Y_{\text{mod}}^{\text{forced}}\in\mathbb{R}^{p}$ to external forcings (such as greenhouse gas emissions, aerosols, solar radiation, or volcanic activity), a crucial step in detecting warming using the fingerprint method (see Hegerl et al., 1996), while ensuring robustness to climate’s Decadal Internal Variability (DIV). To achieve this, we follow Sippel et al. (2021), leveraging multiple climate models from the Climate Model Intercomparison Project (CMIP), both Phase 5 Taylor et al. (2011) and 6 Eyring et al. (2016) (CMIP5 and CMIP6). Specifically, we employ four models (CCSM4, NorCPM1, CESM2, HadCM3) characterised by lower-scale DIV to train our MVA algorithm and validate its robustness using anchor regularisation against models exhibiting higher-scale DIV (CNRM-CM6-1, CNRM-ESM2-1, IPSL-CM6A-LR).
Figure 4: Standard deviation of DIV among training (red) and testing (blue) model members.
Estimators
As a multivariate algorithm, we employ RRRR since both the predictors (temperature fields) and the target (temperature response to external forcing) variables exhibit correlation structures due to spatial autocorrelation. Since both $\alpha$ and $\rho$ in RRRR serve to regularise the regression, we optimise them using cross-validation across the training models. We consider two levels of anchor regularisation: $\gamma=5$ (low regularisation) and $\gamma=100$ to induce high robustness to a potential increase in DIV. We evaluate our results regarding two metrics: $R^{2}$ score (a standard metric when predicting spatial fields) and mean correlation between the anchor (DIV) and the regression residuals (Sippel et al., 2021; Székely et al., 2022).
Data
We selected seven models from the CMIP5 and CMIP6 archives, each containing at least $8$ members from historical simulations, to ensure accurate estimation of the climate response (refer to Tab. 2 for detailed model information). The data preprocessing procedure for each model involves re-gridding surface air temperature data to a regular 5°5  °$\times$5°5  ° grid and computing yearly anomalies by subtracting the mean surface air temperature for the reference period: years $1850-1900$. The forced response $Y^{\text{forced}}_{\text{mod}}$ is obtained using a standard approach (see Deser et al., 2020, and Supplementary details), averaging over all available members in each model. Furthermore, we use DIV as a proxy for multidecadal climate internal variability (see Parsons et al., 2020), achieved by removing the global forced response to global temperature and smoothing it using a $10$-year running mean. This procedure is commonly employed for estimating the multidecadal variability of the climate system (see Sippel et al., 2021; Deser et al., 2020, for further details). Finally, all data are standardised.
Evaluation procedure
We categorise the models into two groups (low and high variability) based on the distribution of the standard deviation of their DIV among the members of each model. As depicted in Fig. 4, a significant proportion of members used for training display a notably smaller magnitude of DIV. Hyperparameters $(\alpha,\rho)$ are selected through a leave-half-models-out cross-validation procedure. Given $k=4$ training models, we randomly select $k/2$ models for training and $k/2$ models for validation, and repeated $B=10$ times. We use $500$ samples from each model to ensure equal weights are given to each training model in the learning algorithm, despite variations in the number of members. For each of the $B$ train/validation splits, we train an RRRR model and select the hyperparameters that yield the highest averaged $R^{2}$ score across the $B$ sampling splits. $R^{2}$ provides a performance measure comparable across regions and is typically used when predicting spatial variables (see e.g. Sippel et al., 2019). Hyperparameter $\alpha$ is selected from $20$ candidates in a logarithmically-spaced sequence ($\lambda\in[1;10^{6}]$), and $\rho$ is selected from $21$ candidates in a linearly-spaced integer sequence ($\rho\in[300;600]$).
Results
For both anchor-regularised RRRR (A-RRRR) with $\gamma=5$ and $\gamma=100$, we observe an improvement in both metrics in the testing set (respectively $0.537$ and $0.533$ of $R^{2}$ for A-RRRR and $0.506$ for unregularised RRRR), while experiencing a slight decrease of performance in the training samples ($0.510$ of $R^{2}$ for unregularised RRRR and respectively $0.500$ and $0.487$ for both A-RRRR). As the latter strongly protects against shifts in DIV, we notice that its mean correlation between residuals and DIV is lower, albeit with a slight decrease in the $R^{2}$ score (see Table 3 in §11). Spatially, we observe (Fig. 5) that anchor-regularised RRRR performs better in the northern hemisphere, particularly in regions where the residuals are highly correlated with DIV and in the northern hemisphere where unregularised RRRR performs poorly (see Figs. 14 and 15 in §11).
Conversely, we observe a decrease in the performance of A-RRRR in regions where unregularised RRRR already performs well or where residuals are anti-correlated with DIV. The latter challenge might be addressed by considering regional proxies of internal variability instead of a global one, which does not always accurately represent local internal variability.
Figure 5: (A) $R^{2}$ score differences between A-RRRR (for $\gamma=5$) and RRRR for testing models. Red values imply that anchor regularisation gives better estimates (larger $R^{2}$ score). (B) Differences of correlation between residuals and DIV ($r$) between anchor-regularised RRRR (for $\gamma=5$) and RRRR for testing models. Red values imply that anchor regularisation gives a lower correlation between residuals and DIV. Regions where A-RRRR perform better are hatched with dots.
These results suggest the potential of using anchor-regularised multi-output algorithms in D&A studies to detect and attribute local responses to external forcings, a fundamental problem in climate science.
7 Conclusion
In this study, we extend the causal framework of anchor regression proposed by Rothenhäusler et al. (2018), demonstrating the versatility of their regularisation approach across a wide range of MVA algorithms. Given the significant challenge of generalising models to OOD data in machine learning, we advocate for integrating anchor regularisation into a broader class of MVA algorithms, including RRR, OPLS, or PLS regression, particularly when prior knowledge is available about the reasons for distribution shifts. Moreover, we highlight that anchor regularisation offers an interesting trade-off between prediction performance and invariance to distribution shifts, addressing concerns regarding over-conservativeness in some cases. Additionally, our application of anchor regularisation in real-world scenarios underscores its practical utility. Future theoretical advancements will entail extending the formulation nonlinear cases using kernel methods Po-Chedley et al. (2022). Furthermore, an intriguing avenue for exploration is understanding how statistical learning algorithms incompatible with anchor regularisation behave when this regularisation is applied. On the application front, we aim to investigate how anchor regularisation, in conjunction with various MVA algorithms, can be leveraged to detect forced responses to more complex climate variables (e.g., precipitation or temperature and their extremes) and attribute them to external forcing sources such as greenhouse gas or aerosol emissions (anthropogenic) or natural phenomena like solar radiation and volcanic activity. We anticipate a broad development and application of the presented and future anchor-regularised methods across various fields of science.
Impact statement
This paper contributes to Machine Learning by showcasing the application of anchor regularisation in both general tasks in multivariate data analysis and statistics and in the specific context of Detection and Attribution (D&A) studies in climate science. The broader impact of this work lies in its potential to enhance the robustness and reliability of algorithms across various domains, including those with societal, economic and environmental implications. Specifically, in climate science, where accurate detection and attribution of climate change impacts are crucial for informed decision-making, our findings demonstrate how anchor regularisation can improve the reliability of predictive models. By addressing distributional shifts and enhancing model robustness, our approach can potentially strengthen the scientific basis for policy-making and adaptation strategies while contributing to the broader understanding and mitigation of climate change impacts.